Science Report – November 14th

Science Report 14 November 2019
SOL: 4
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: Gypsum samples retrieved were cleaned and weighed before
commencing grinding prior to experimentation. These samples were joined by
micrometeorite samples which were processed and packaged for shipment.
Green Hab: Green Hab operations have continued. See Green Hab report.
RAM Operations: N/A
EVA: The fourth EVA for science saw us returning to the MM5 (518462E 4248949N) site
where 30 micrometeorite samples were collected. 40% of the grid has now been sampled.
See EVA Report.
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin
monitoring vests. The simulated emergency retrieval of the engineer has been postponed,
subject to approval, to 15Nov2019 at the conclusion of his engineering EVA.
Additional activities: Hab operations were disrupted due to ongoing construction.

Copyright © The Mars Society. All rights reserved. | Main Site